# The cohomological Kudla conjecture for unitary Shimura varieties

### v1

@@ 1. Introduction
Our study features a natural class of compactifications of Shimura varieties called toroidal compactifications, first introduced in [AMRT10]. In the case of orthogonal Shimura varieties, the following conjecture was originally posed by Kudla [Kud04, Problem 3] and then reformulated more recently by Bruinier-Rosu-Zemel [HRZ24].
@@ 1.1. Main results
Let $k$ be an imaginary quadratic field with ring of integers $\mathcal{O}_{k}$ and let $(V,h)$ be a Hermitian vector space over $k$ of signature $(n+1,1)$. Let $L\subset V$ be an $\mathcal{O}_{k}$-lattice, i.e., $h$ is $\mathcal{O}_{k}$-valued on $L$, and $L\otimes_{\mathcal{O}_{k}}k=V$. One can attach to this data a unitary Shimura variety $X_{\Gamma}$ whose construction is recalled in Section 3. Let $11\leq g\leq n+$ 1, $\underline{\nu}\in(L^{\vee}/L)^{g}$ and let $N\in\mathrm{Herm}_{g}(k)_{\geq 0}$ be a Hermitian semi-positive matrix. For each $(\underline{\nu},N)$, there is a special cycle ${\mathcal{Z}}(\underline{\nu},N)\hookrightarrow X_{\Gamma}$ of codimension $g$. Let
@@
$\Phi_{L}^{g}(\tau)=\sum_{\underset{\underline{\nu}\in(L^{\vee}/L)^{g}}{N\in\mathrm{Herm}_{g}(k)_{\geq 0}}}[{\mathcal{Z}}(\underline{\nu},N)]q^{N}\mathfrak{e}_{\underline{\nu}}\in\mathrm{CH}^{g}(X_{\Gamma})\otimes\mathbb{C}[(L^{\vee}/L)^{g}]\llbracket q\rrbracket\leavevmode\nobreak\ ,$
where $\tau\in{\mathcal{H}}_{g}$ is an element of the Hermitian upper half-space and $q^{N}=e^{2i\pi\mathrm{Tr}(\tau N)}$.
There is a unique toroidal compactification $X_{\Gamma}\subset X_{\Gamma}^{\mathrm{tor}}$, and we denote by $\overline{{\mathcal{Z}}}(\underline{\nu},N)$ the Zariski closure of the special cycle ${\mathcal{Z}}(\underline{\nu},N)$ in $X^{\mathrm{tor}}_{\Sigma}$. Our first main result is the following.
@@
Assume that $g\leq\frac{n}{2}$. There exists a correction $\widetilde{{\mathcal{Z}}}(\underline{\nu},N)\subset X^{\mathrm{tor}}_{\Sigma}$ to each $\overline{{\mathcal{Z}}}(\underline{\nu},N)$, by a cycle supported in the boundary $X^{\mathrm{tor}}_{\Sigma}\setminus X_{\Sigma}$, such that the generating series:
$\sum_{\underset{\underline{\nu}\in(L^{\vee}/L)^{g}}{N\in\mathrm{Herm}_{g}(k)_{\geq 0}}}[\widetilde{{\mathcal{Z}}}(\underline{\nu},N)]q^{N}\mathfrak{e}_{\underline{\nu}}\in\mathrm{H}^{2g}\left(X^{\mathrm{tor}}_{\Gamma}\right)\otimes\mathbb{C}[(L^{\vee}/L)^{g}]\llbracket q\rrbracket\leavevmode\nobreak\ .$
@@
To give a more explicit form to the corrections and completions in Theorems 1.2 and 1.3, we need to recall the structure of the toroidal compactification $X_{\Gamma}^{\mathrm{tor}}$. For each equivalence class of primitive isotropic line $\mathfrak{J}\subset L$ under the action of $\Gamma$, we have an associated boundary divisor ${\mathcal{B}}_{\mathfrak{J}}$ which is isomorphic to $E_{M}$, where $M=\mathfrak{J}^{\bot}/\mathfrak{J}$, and $E=\mathbb{C}/{\mathcal{O}}_{k}$ is a CM elliptic curve. Let $\iota_{\mathfrak{J}}:{\mathcal{B}}_{\mathfrak{J}}\hookrightarrow X_{\Gamma}^{\mathrm{tor}}$ denote the inclusion and let
$\iota_{{\mathfrak{J}},*}:H^{*}({\mathcal{B}}_{\mathfrak{J}},\mathbb{Q})\rightarrow H^{*+2}(X_{\Gamma}^{\mathrm{tor}},\mathbb{Q})\leavevmode\nobreak\ ,$
@@
$\mathrm{H}^{2g}({\mathcal{B}}_{\mathfrak{J}},\mathbb{Q})\simeq\bigoplus_{0\leq\ell\leq g}{\mathcal{L}}^{g-\ell}\mathrm{H}^{2\ell}_{\mathrm{prim}}(\mathcal{B}_{\mathfrak{J}},\mathbb{Q})\leavevmode\nobreak\ .$
@@
$\mathrm{Hdg}^{g}({\mathcal{B}}_{\mathfrak{J}})\simeq\bigoplus_{0\leq\ell\leq g}{\mathcal{L}}^{g-\ell}\mathrm{Hdg}^{\ell}_{\mathrm{prim}}(\mathcal{B}_{\mathfrak{J}})\leavevmode\nobreak\ .$
For each $0\leq\ell\leq\frac{n}{2}$, consider a basis $(W_{i}^{\ell})_{i\in I_{\ell}}$ of the primitive cohomology $\mathrm{Hdg}^{\ell,\ell}_{\mathrm{prim}}(\mathcal{B}_{\mathfrak{J}},\mathbb{Q})$. As the Hodge conjecture holds for ${\mathcal{B}}_{\mathfrak{J}}$, the $W_{i}^{\ell}$ are classes of algebraic cycles; see 3.6. To each primitive cycle $W^{\ell}_{i}$, we associate a harmonic polynomial [^1] Polynomial in the real and imaginary parts of the $\lambda_{i}$’s. $P^{\ell}_{i}$ in $(\lambda_{1},\ldots,\lambda_{\ell})\in(L_{\mathbb{R}})^{\ell}$ that satisfies the following equivariance property for $A\in\mathrm{GL}_{\ell}(\mathbb{C})$:
$P\left((\lambda_{1},\ldots,\lambda_{\ell})\cdot A\right)=|\det(A)|^{2}\cdot P(\lambda_{1},\ldots,\lambda_{\ell})\leavevmode\nobreak\ .$
The space of all polynomials that satisfy this homogeneity condition is in fact a finite-dimensional vector space denoted by ${\mathcal{F}}_{n,\ell}$, and the direct sum
@@
admits an $\mathfrak{sl}_{2}$-action introduced in Section 2.2, the primitive elements of which are exactly the harmonic polynomials, i.e., the span of the $P_{i}^{\ell}$. Let $(E,F,H)$ be the corresponding $\mathfrak{sl}_{2}$ triple. Our main theorem is the following.
@@
[…]
$[\widetilde{{\mathcal{Z}}}(\underline{\nu},N)]=[{\mathcal{Z}}(\underline{\nu},N)]+\sum_{{\mathfrak{J}}/\sim}\sum_{\underset{i\in I_{\ell}}{0\leq\ell\leq g-1}}\frac{r_{\mathfrak{J}}}{d_{k}}P^{\ell}_{i}(\underline{\nu},N)\,\mathfrak{\iota}_{{\mathfrak{J}},*}[W^{\ell}_{i}\cup\mathcal{L}^{g-\ell-1}]\leavevmode\nobreak\ ,$
@@
$P^{\ell}_{i}(\underline{\nu},N):=\sum_{\begin{subarray}{c}(\lambda_{1},\ldots,\lambda_{g})\in(M^{\vee})^{g},\\ {\mathrm{Gram}}(\underline{\lambda})=N,\\ [\underline{\lambda}]=\uparrow^{M}_{L}(\underline{\nu})\\ \end{subarray}}(E^{g-\ell}P^{\ell,\ell}_{i})(\lambda_{1},\ldots,\lambda_{g})\leavevmode\nobreak\ ,$
$E$ is the raising operator on the $\mathfrak{sl}_{2}$-module ${\mathcal{F}}_{n,\bullet}$, $-d_{k}$ is the discriminant of $k$, and $r_{\mathfrak{J}}$ is a certain rational number associated to ${\mathfrak{J}}$, see Section 3.3.
Let $\vartheta_{M,P}$ denotes the weighted theta series of the homogeneous polynomial $P$, then we get the following corollary.
@@
The corrected sum:
$\Phi_{L}^{g}+\sum_{{\mathfrak{J}}/\sim}\sum_{\underset{i\in I_{\ell}}{0\leq\ell\leq g-1}}\frac{r_{\mathfrak{J}}}{d_{k}}\cdot\uparrow_{M}^{L}(\vartheta_{M,{P^{\ell}_{i}}})\otimes[W_{i}^{\ell}\cup\mathcal{L}^{g-\ell-1}]$
@@
The theta series $\vartheta_{M,{P^{\ell}_{i}}}$ admit non-holomorphic completions which transform like modular forms; see Section 2.4, whence the following corollary.
@@
The generating series $\Phi_{L}^{g}$ is a Hermitian quasi-modular form with values in the cohomology of $X^{\mathrm{tor}}_{\Gamma}.$
We refer to Definition 5.1 for an explicit expression of the non-holomorphic completions of the special cycles.
@@
The statements of Theorem 1.3 and Theorem 1.4 still hold in the case of $\mathrm{U}(n+1,1)$, with values in the Chow groups of all codimensions.
@@ 1.2. Previous and related work
The modularity of generating series of special cycles on Shimura varieties first appeared in the work of Hirzerbruch and Zagier in the context of Hilbert modular surfaces [HZ76]. Subsequent work of Kudla and Millson [KM90] showed that generating series of cohomology classes of cycles of arbitrary codimension in locally symmetric spaces associated to orthogonal (resp. unitary) groups give rise to holomorphic Siegel (resp. Hermitian) modular forms. If the locally symmetric space is associated to an orthogonal group of signature $(n,2)$ or a unitary group, then it is a Shimura variety, and stronger results are known. Borcherds [Bor99, Bor98] proved that the generating series of special divisors in orthogonal Shimura varieties is a modular form with values in the Picard group, and Bruinier–Westerholt-Raum [BWR15], based on previous work of Zhang [Zha09], proved that the generating series of higher codimension cycles is a Siegel modular form, also valued in the Chow group. Similar results are also known in the case of unitary Shimura varieties [Xia22, Liu11, YZZ09], and we refer to the introduction of [HRZ24] for a more detailed background.
On the other hand, the behavior of the generating series of compactified special cycles in a toroidal compactification is more mysterious. For orthogonal Shimura varieties, the case of divisors was first solved by Bruinier-Zemel [BZ21]; see also subsequent work [EGT23, Gar23]. For zero-cycles, Bruinier-Rosu-Zemel [HRZ24] have recently proven Conjecture 1.1 in the orthogonal case and under some assumptions in the unitary case, see [HRZ24, Theorems 1.2, 1.3, 1.4]. Therefore, Theorem 1.4 is the first general result showing modularity of compactified cycles of arbitrary codimension, and it also provides an explicit non-holomorphic completion.
@@ 1.3. Strategy of proof
The proof of Theorem 1.4 relies on several ingredients:
@@
  • The boundary divisors can be explicitly understood as finite group quotients of $E_{M}$ where $E$ is the CM elliptic curve $\mathbb{C}/{\mathcal{O}}_{k}$, and $M$ is a projective $\mathcal{O}_{k}$ module with positive definite Hermitian pairing. We construct Poincaré dual harmonic forms to the special cycles in each boundary divisor, which appear to be an instance of theta lifting. A modularity criterion of Shimura allows us to explicitly construct non-holomorphic completions.
  • To construct the corrections to the special cycles, we first develop a theory of Hermitian quasi-modular forms using theta series weighted by certain homogeneous polynomials. We show that the space of such polynomials admits an $\mathfrak{sl}_{2}$-action and the special cycles intertwine this action with the $\mathfrak{sl}_{2}$ action on the cohomology of the boundary. This compatibility leads to a proof of Theorem 1.4.

### v2

@@ 1. Introduction
Our study features a natural class of compactifications of Shimura varieties called toroidal compactifications, first introduced in [AMRT10]. In the case of orthogonal Shimura varieties, the following conjecture was originally posed by Kudla [Kud04, Problem 3], see also the recent reformulation of Bruinier-Rosu-Zemel [BRZ24].
@@ 1.1. Main results
Let $k$ be an imaginary quadratic field with ring of integers $\mathcal{O}_{k}$ and let $(V,h)$ be a Hermitian vector space over $k$ of signature $(n+1,1)$. Let $L\subset V$ be an $\mathcal{O}_{k}$-lattice, i.e., $h$ is $\mathcal{O}_{k}$-valued on $L$, and $L\otimes_{\mathcal{O}_{k}}k=V$. One can attach to this data a unitary Shimura variety $X_{\Gamma}$ whose construction is recalled in Section˜3. Let $1\leq g\leq n+1$, $\underline{\nu}\in(L^{\vee}/L)^{g}$ and let $N\in\mathrm{Herm}_{g}(k)_{\geq 0}$ be a Hermitian semi-positive matrix. For each $(\underline{\nu},N)$, there is a special cycle ${\mathcal{Z}}(\underline{\nu},N)\hookrightarrow X_{\Gamma}$ of codimension $g$. Let
@@
(1.1) $\sum_{\underset{\underline{\nu}\in(L^{\vee}/L)^{g}}{N\in\mathrm{Herm}_{g}(k)_{\geq 0}}}[{\mathcal{Z}}(\underline{\nu},N)]q^{N}\mathfrak{e}_{\underline{\nu}}\in\mathrm{CH}^{g}(X_{\Gamma})\otimes\mathbb{C}[(L^{\vee}/L)^{g}]\llbracket q\rrbracket\penalty 10000\ ,$
where $\tau\in{\mathcal{H}}_{g}$ is an element of the Hermitian upper half-space and $q^{N}=e^{2i\pi\mathrm{Tr}(\tau N)}$. In this setting, with values in $\mathrm{CH}^{g}(X_{\Gamma})$, the convergence of the series above is still open; the analogous issue was settled in the orthogonal case by Bruinier-Westerholt-Raum [BWR15]. For the rest of this paper, we will work with generating series valued in cohomology and we denote by $\Phi_{L}^{g}$ the series obtained by taking cohomology classes in Equation˜1.1.
There is a unique toroidal compactification $X_{\Gamma}\subset X_{\Gamma}^{\mathrm{tor}}$, and we denote by $\overline{{\mathcal{Z}}}(\underline{\nu},N)$ the Zariski closure of the special cycle ${\mathcal{Z}}(\underline{\nu},N)$ in $X^{\mathrm{tor}}_{\Gamma}$. Our first main result is the following.
@@
Assume that $g\leq\frac{n}{2}$. There exists a correction $\widetilde{{\mathcal{Z}}}(\underline{\nu},N)\subset X^{\mathrm{tor}}_{\Gamma}$ to each $\overline{{\mathcal{Z}}}(\underline{\nu},N)$, by a cycle supported in the boundary $X^{\mathrm{tor}}_{\Gamma}\setminus X_{\Gamma}$, such that the generating series:
$\widetilde{\Phi}_{L}^{g}(\tau)=\sum_{\underset{\underline{\nu}\in(L^{\vee}/L)^{g}}{N\in\mathrm{Herm}_{g}(k)_{\geq 0}}}[\widetilde{{\mathcal{Z}}}(\underline{\nu},N)]q^{N}\mathfrak{e}_{\underline{\nu}}\in\mathrm{H}^{2g}\left(X^{\mathrm{tor}}_{\Gamma}\right)\otimes\mathbb{C}[(L^{\vee}/L)^{g}]\llbracket q\rrbracket\penalty 10000\ .$
@@
To give a more explicit form to the corrections and completions in Theorems˜1.2 and 1.3, we need to recall the structure of the toroidal compactification $X_{\Gamma}^{\mathrm{tor}}$. For each equivalence class of primitive isotropic line $\mathfrak{J}\subset L$ under the action of $\Gamma$, we have an associated boundary divisor ${\mathcal{B}}_{\mathfrak{J}}$ which is isomorphic to the Serre tensor product $E_{M}=M\otimes E$, where $M=\mathfrak{J}^{\bot}/\mathfrak{J}$, and $E=\mathbb{C}/{\mathcal{O}}_{k}$ is a CM elliptic curve. Let $\iota_{\mathfrak{J}}:{\mathcal{B}}_{\mathfrak{J}}\hookrightarrow X_{\Gamma}^{\mathrm{tor}}$ denote the inclusion and let
$\iota_{{\mathfrak{J}},*}:\mathrm{H}^{*}({\mathcal{B}}_{\mathfrak{J}},\mathbb{Q})\rightarrow\mathrm{H}^{*+2}(X_{\Gamma}^{\mathrm{tor}},\mathbb{Q})\penalty 10000\ ,$
@@
$\mathrm{H}^{2g}({\mathcal{B}}_{\mathfrak{J}},\mathbb{Q})\simeq\bigoplus_{0\leq\ell\leq g}{\mathcal{L}}^{g-\ell}\mathrm{H}^{2\ell}_{\mathrm{prim}}(\mathcal{B}_{\mathfrak{J}},\mathbb{Q})\penalty 10000\ .$
@@
$\mathrm{Hdg}^{g}({\mathcal{B}}_{\mathfrak{J}})\simeq\bigoplus_{0\leq\ell\leq g}{\mathcal{L}}^{g-\ell}\mathrm{Hdg}^{\ell}_{\mathrm{prim}}(\mathcal{B}_{\mathfrak{J}})\penalty 10000\ .$
For each $0\leq\ell\leq\frac{n}{2}$, consider a basis $(W_{i}^{\ell})_{i\in I_{\ell}}$ of the primitive cohomology $\mathrm{Hdg}^{\ell}_{\mathrm{prim}}(\mathcal{B}_{\mathfrak{J}})$. As the Hodge conjecture holds for ${\mathcal{B}}_{\mathfrak{J}}$, the $W_{i}^{\ell}$ are classes of algebraic cycles; see ˜3.6. To each primitive cycle $W^{\ell}_{i}$, we associate in Section 4.3 a harmonic polynomial [^1] Polynomial in the real and imaginary parts of the $\lambda_{i}$’s. $P^{\ell,\ell}_{i}$ in $(\lambda_{1},\ldots,\lambda_{\ell})\in(L_{\mathbb{R}})^{\ell}$ that satisfies the following equivariance property for $A\in\mathrm{GL}_{\ell}(\mathbb{C})$:
$P\left((\lambda_{1},\ldots,\lambda_{\ell})\cdot A\right)=|\det(A)|^{2}\cdot P(\lambda_{1},\ldots,\lambda_{\ell})\penalty 10000\ .$
The space of all polynomials that satisfy the previous homogeneity condition is in fact a finite-dimensional vector space denoted by ${\mathcal{F}}_{n,\ell}$, and the direct sum
@@
admits an $\mathfrak{sl}_{2}$-action introduced in Section˜2.3, the primitive elements of which are exactly the harmonic polynomials, i.e., the span of the $(P_{i}^{\ell,\ell})_{\ell,i\in I_{\ell}}$. Let $(\Lambda,F,H)$ be the corresponding $\mathfrak{sl}_{2}$-triple. Our main theorem is the following.
@@
Let $g\leq\frac{n}{2}$, let $\underline{\nu}\in(L^{\vee}/L)^{g}$, and let $N\in\mathrm{Herm}_{g}(k)_{\geq 0}$ be a Hermitian semi-positive matrix. In the cohomology group $H^{2g}(X_{\Gamma}^{\mathrm{tor}},\mathbb{Q})$, the correction of the cycle $\overline{{\mathcal{Z}}}(\underline{\nu},N)$ is trivial if $p_{L}^{M}(\underline{\nu})=0$ or if $N$ is not positive definite  [^2] This operator between Weil representations is defined in Section 2.2. and otherwise, it is given by the formula:
$\widetilde{{\mathcal{Z}}}(\underline{\nu},N)]=[\overline{{\mathcal{Z}}}(\underline{\nu},N)]+\sum_{{\mathfrak{J}}/\sim}\sum_{\underset{i\in I_{\ell}}{0\leq\ell\leq g-1}}\frac{r_{\mathfrak{J}}}{d_{k}}P^{\ell}_{i}(\underline{\nu},N)\,\mathfrak{\iota}_{{\mathfrak{J}},*}[W^{\ell}_{i}\cup\mathcal{L}^{g-\ell-1}]\penalty 10000\ ,$
@@
$P^{\ell}_{i}(\underline{\nu},N):=\sum_{\begin{subarray}{c}(\lambda_{1},\ldots,\lambda_{g})\in(M^{\vee})^{g},\\ {\mathrm{Gram}}(\underline{\lambda})=N,\\ [\underline{\lambda}]=p^{M}_{L}(\underline{\nu})\\ \end{subarray}}(\Lambda^{g-\ell}P^{\ell,\ell}_{i})(\lambda_{1},\ldots,\lambda_{g})\penalty 10000\ ,$
$\Lambda$ is the raising operator on the $\mathfrak{sl}_{2}$-module ${\mathcal{F}}_{n,\bullet}$, $d_{k}$ is the discriminant of $k$, and $r_{\mathfrak{J}}$ is a certain rational number associated to $\mathfrak{J}$}, see Section˜3.3.
Let $\vartheta_{M,P}(\tau)$ denote the weighted theta series of the homogeneous polynomial $P$, then we get the following corollary.
@@
The cohomology class of the corrected sum:
$\widetilde{\Phi}^{g}_{L}(\tau)=\overline{\Phi}_{L}^{g}(\tau)+\sum_{{\mathfrak{J}}/\sim}\sum_{\underset{i\in I_{\ell}}{0\leq\ell\leq g-1}}\frac{r_{\mathfrak{J}}}{d_{k}}\cdot p_{M}^{L}(\vartheta_{M,{P^{\ell}_{i}}}(\tau))\otimes\mathfrak{\iota}_{{\mathfrak{J}},*}[W_{i}^{\ell}\cup\mathcal{L}^{g-\ell-1}]$
@@
The theta series $\vartheta_{M,{P^{\ell}_{i}}}(\tau)$ admit non-holomorphic completions which transform like modular forms; see Section˜2.5, whence the following corollary.
@@
The generating series $\overline{\Phi}_{L}^{g}$ is a Hermitian quasi-modular form with values in the cohomology of $X^{\mathrm{tor}}_{\Gamma}.$
We refer to Definition˜5.1 for an explicit expression of the non-holomorphic completions of the special cycles.
@@
The statements of Theorem˜1.3 and Theorem˜1.4 still hold in the case of $\mathrm{U}(n+1,1)$, with values in the Chow groups of all codimensions.
@@ 1.2. Previous and related work
The modularity of generating series of special cycles on Shimura varieties first appeared in the work of Hirzebruch and Zagier in the context of Hilbert modular surfaces [HZ76]. Subsequent work of Kudla and Millson [KM90] showed that generating series of cohomology classes of cycles of arbitrary codimension in locally symmetric spaces associated to orthogonal (resp. unitary) groups give rise to holomorphic Siegel (resp. Hermitian) modular forms. If the locally symmetric space is associated to an orthogonal group of signature $(n,2)$ or a unitary group, then it is a Shimura variety, and stronger results are known. Borcherds [Bor99, Bor98] proved that the generating series of special divisors in orthogonal Shimura varieties is a modular form with values in the Picard group, and Bruinier–Westerholt-Raum [BWR15], based on previous work of Zhang [Zha09], proved that the generating series of higher codimension cycles is a Siegel modular form, also valued in the Chow group. Similar results are also known in the case of unitary Shimura varieties [Xia22, Liu11, YZZ09], and we refer to the introduction of [BRZ24] for a more detailed background.
On the other hand, the behavior of the generating series of compactified special cycles in a toroidal compactification is more mysterious. For orthogonal Shimura varieties, the case of divisors was first solved by Bruinier-Zemel [BZ21]; see also subsequent work [EGT23, Gar23] and for divisors in unitary Shimura varieties of signature $(n,1)$ it is a consequence of [BHK+20], see also the work of Cogdell [Cog85]. For zero-cycles, Bruinier-Rosu-Zemel [BRZ24] have recently proven Conjecture 1.1 in the orthogonal case and under some assumptions in the unitary case, see [BRZ24, Theorems 1.2, 1.3, 1.4]. Therefore, Theorem˜1.4 is the first general result showing modularity of compactified cycles of arbitrary codimension, and it also provides an explicit non-holomorphic completion.
@@ 1.3. Strategy of proof
The proof of Theorem˜1.4 relies on several ingredients:
@@
  • The boundary divisors can be explicitly understood as finite group quotients of $E_{M}$ where $E$ is the CM elliptic curve $\mathbb{C}/{\mathcal{O}}_{k}$, and $M$ is a projective $\mathcal{O}_{k}$ module with positive definite Hermitian pairing. We construct Poincaré dual harmonic forms to the special cycles in each boundary divisor, which appear to be an instance of theta lifting. A modularity criterion allows us to explicitly construct non-holomorphic completions.
  • To construct the corrections to the special cycles, we first develop a theory of Hermitian quasi-modular forms using theta series weighted by certain homogeneous polynomials. We show that the space of such polynomials admits an $\mathfrak{sl}_{2}$-action and the special cycles intertwine this action with the $\mathfrak{sl}_{2}$-action on the cohomology of the boundary. This compatibility leads to a proof of Theorem˜1.4.